# Deep Common Feature Mining for Efficient Video Semantic Segmentation

### v1

@@ 4.1 Datasets and evaluation metrics
[…]
@@
Evaluation metrics. We evaluate our approach using mean Intersection over Union (mIoU) and frequency-weighted Intersection over Union (wIoU) metrics on annotated frames. Additionally, we measure inference speed in frames per second (FPS). For assessing temporal consistency, we use mean Video Consistency (mVC) [28].
4.2 Implementation details
During training, data augmentation techniques such as random rotation, random resizing, random cropping, color jitter, and random horizontal flipping are implemented. The crop size is set to $480\times 480$ for VSPW and $768\times 1536$ pixels for Cityscapes. At the inference stage, the entire frame is fed into the model on a single Tesla V100 GPU, avoiding the sliding window input. Frame sizes are resized to $480\times 853$ for VSPW and $512\times 1024$ for Cityscapes when using MiT as the backbone. The default interval $K$ between keyframes is set to 2. Additional experimental details are available in the supplementary materials.
Methods Backbone mIoU $\uparrow$ FPS $\uparrow$
@@ 4.2 Implementation details
Table 2: Performance comparison in terms of accuracy (mIoU) and inference speed (FPS) on the Cityscapes [10] validation dataset. Methods, excluding ours, are arranged based on mIoU. DCFM* indicates a variant of our model with a different input resolution ($640\times 1280$). Results are visualized in Figure LABEL:fig:sota-cs.
4.3 Comparison with previous approaches
We comprehensively compare our method with the existing ones on VSPW [28]. Results in Table 1 highlight the notable performance of our models, particularly DCFM-B2 and DCFM-B5. Compared to the image baseline (SegFormer [43]), DCFM-B2 and DCFM-B5 improve FPS from 44.1 and 13.9 to 67.6 and 25.4 (speed up 1.5 times and 1.8 times), respectively, while maintaining almost the same accuracy with the keyframe interval $K$ = 2. In comparison to MRCFA-B5, the current state-of-the-art method on VSPW in terms of accuracy, DCFM-B5 ($K$ = 2) exhibits a slight decrease of -1.7% in mIoU, but it achieves around fivefold increase in FPS (from 5.3 to 25.4). By adjusting $K$ to 6 during inference, our DCFM-B5 model achieves a significantly faster speed. Despite the slight decrease in mIoU, it still outperforms MRCFA-B2 and CFFM-B2 in terms of both accuracy and inference speed, despite having a larger parameter count. The results visualized in Figure LABEL:fig:sota-vs further demonstrate superior temporal consistency (mVC) and a more favorable trade-off between accuracy and efficiency achieved by the proposed approach.
@@ 4.3 Comparison with previous approaches
We extend our model comparison to the Cityscapes [10], detailed in Table 2 and Figure LABEL:fig:sota-cs. DCFM-50 outperforms contemporary feature aggregation methods like TDNet, which prioritizes high accuracy while utilizing multiple distributed networks for enhanced inference speed simultaneously. Notably, our method achieves a superior mIoU while maintaining a higher inference speed. In comparison to the method with the highest mIoU score, LMA [30], our DCFM-101 exhibits a 0.5% lower mIoU but is 4.3 times faster. When compared with other keyframe-based techniques aimed at accelerating segmentation models, we surpass LVS [31] and DCNet [1] in both accuracy (77.2% mIoU) and speed (13.2 FPS) with a smaller backbone (ResNet-50 [14]). These two methods also leverage different layers of the backbone network for feature reuse like ours, while our approach stands out by decoupling keyframe information into the common and independent parts. Among the methods using MiT [43] as the backbone, DCFM is both faster and more accurate than the SegFormer [43] baseline. Compared with CFFM [36] and MRCFA [36], the most accurate methods on VSS, DCFM achieves 0.7% less mIoU but about twice as fast.
Results on both datasets affirm the effectiveness of DCFM. Further insights into the advantage of such feature decoupling are detailed in Section 4.4.
4.4 Ablation study
Effectiveness of feature decomposing. By decomposing features into common and independent representations, our approach distinguishes itself from other keyframe-based methods. In Table 3, we perform ablation study by omitting one of the feature components in keyframes or non-key frames during both training and testing. As shown in the first row, excluding the shallow features $F^{key}_{in}$ of the keyframe causes our framework to revert to the structure of previous feature propagation-based methods, as depicted in Figure LABEL:fig:moti_b. The observed significant decrease in accuracy suggests a misalignment between the keyframes and the corresponding non-keyframes. However, instead of applying a re-calibration operation on the utilized feature to address this issue, our framework mitigates misalignments by mining common information across frames. Visualization of segmentation results are available in the supplementary material. Comparing the results of the ablation model in the second row with those obtained by the full model, it is evident that the learned common feature is highly informative, containing most of the necessary information for the non-key frames. However, utilizing the common feature $F^{key}_{co}$ provided by the keyframe alone on non-key frames is suboptimal. This is primarily because it neglects the quickly changing frame-specific dynamics, and such complementary information is well captured by the independent feature in our full model.
@@ 4.4 Ablation study
42.0 41.8 41.9
41.8 40.8 41.3
43.9 43.6 43.7
[…]
$L_{C}$ mIoU $\text{mVC}_{8}$ $\text{mVC}_{16}$
w/o 42.3 87.3 82.5
w/ 43.7 87.7 83.2
Table 4: Quantitative ablation study on the consistency loss function, achieved on the VSPW [28] validation dataset.
Effectiveness of consistency loss. Table 4 illustrates that the inclusion of $L_{C}$ significantly enhances segmentation accuracy and temporal consistency. Figure 4 visually evaluates the impact by calculating cosine similarity between each pixel and its surrounding eight pixels in the feature space (output of FFM), indicating that the learned features are less susceptible to rapidly changing details in areas with the same semantic category when applying $L_{C}$.
@@
Table 5: Comparison of the inference speed on VSPW [28].
Figure 5: Results of adaptive keyframe scheduling on VSPW [28].
@@
Analysis on inference speed. As shown in Table 5, for a keyframe, DCFM only uses an additional lightweight FFM to integrate features before the header, the inference time increases by only 1.1ms compared with the SegFormer-B5 [43] baseline. Regarding a non-key frame, the computing capacity of the shallow network has only 1.48 GFLOPS, thus DCFM achieves a very fast speed of 5.8ms (172 FPS) by reusing the keyframe common feature, constituting only 8% of the keyframe inference time.
[…]
5 Conclusion
@@
This study addresses challenges in efficient video semantic segmentation. The proposed DCFM method aims to enhance accuracy and temporal consistency while overcoming computational bottlenecks and feature propagation in existing approaches. It decomposes the intricate features into two components, i.e., common and independent representations, where the former can be directly reutilized by neighboring frames without additional re-calibration, contributing to an optimized trade-off between accuracy and speed. A symmetric training strategy and a self-supervised consistency loss are introduced, enhancing the ability to handle sparsely annotated data, offering potential applications in other video analysis tasks.

### v2

@@ IV-A Datasets and evaluation metrics
Cityscapes. Cityscapes [26] consists of 5,000 videos, divided into training (2,975 videos), validation (500 videos), and test sets (1,525 videos). The dataset includes 19 semantic categories, and each video contains 30 frames at a frame rate of 17 FPS. The frame resolution is $1024\times 2048$ pixels, with fine annotations available only on the 20th frame.
CamVid. CamVid [69] comprises 4 videos with a frame rate of 30 FPS. Annotations are provided for every 30th frame, covering 11 semantic categories. The dataset includes a total of 701 annotated frames, divided into 367 for training, 101 for validation, and 233 for testing.
@@
Evaluation metrics. We evaluate our approach using mean Intersection over Union (mIoU) [27] and frequency-weighted Intersection over Union (wIoU) [27] metrics on annotated frames. Additionally, we measure inference speed in frames per second (FPS). For assessing temporal consistency, we use mean Video Consistency (mVC) [14]. For a video containing $n$ frames $\{x_{1},x_{2},...,x_{n}\}$, $\text{VC}_{l}$ represents the consistency of the segmentation results calculated for every $l$ adjacent frames and then averaged:
$\text{VC}_{l}=\frac{1}{n-l+1}\sum_{i=1}^{n-l+1}\frac{(\cap_{i}^{i+l-1}y_{i})\cap(\cap_{i}^{i+l-1}\hat{y}_{i})}{\cap_{i}^{i+l-1}y_{i}},$ (13)
where $y_{i}$ and $\hat{y}_{i}$ represent the label and the segmentation result of the $i^{th}$ frame, respectively. Then, mVCl is obtained by averaging the VCl of each video.
IV-B Implementation details
We provide details of our experimental setup in Table I. When $C_{F}=C_{out}$, we eliminate the convolutional layer depicted in Figure 3a, as $C_{out}$ is sufficiently small in MixTransformer (MiT) [2]. During training, we employ mini-batch stochastic gradient descent (SGD) and AdamW as the optimizers for the backbone networks ResNet [1] and MiT [2], following the practices of PSPNet [33] and SegFormer [2]. We implement data augmentation techniques such as random rotation at a slight angle, random resizing, random cropping, color jitter, and random horizontal flipping. The crop size is set to $480\times 480$, $768\times 1536$ and $640\times 640$ pixels for VSPW [14], Cityscapes [26] and CamVid [69], respectively. The models are trained on 4 GPUs with SyncBN [70], utilizing a batch size of 16, 8 and 8 for the VSPW, Cityscapes and CamVid datasets, respectively. For models using ResNet [1] as the backbone network, we adopt a larger crop size due to ResNet being more sensitive to input size compared to ViT models.
At the inference stage, the entire frame is fed into the model on a single Tesla V100 GPU, avoiding the sliding window input. Frame sizes are resized to $480\times 853$ for VSPW and $512\times 1024$ for Cityscapes when using MiT as the backbone. The default interval $K$ between keyframes is set to 2. In the case of inference on VSPW [14], which is annotated frame-by-frame, we designate the $[(i\mod K)+1]^{th}$ frame as the first keyframe of the $i^{th}$ video. When conducting tests on Cityscapes [26] and CamVid [69], we sequentially set the first keyframe of each video to frames $1$, $2$, up to $K$ and then average the results, given that only the $20^{th}$ frame is annotated. In experiments involving adaptive keyframe scheduling (AKS), we designate the $1^{st}$ frame as the initial keyframe to ensure result stability and facilitate comparisons.
IV-C Comparison with previous approaches
We comprehensively compare our method with the existing ones on VSPW [14]. Results in Table II highlight the notable performance of our models, particularly DCFM-B2 and DCFM-B5. Compared to the image baseline (SegFormer [2]), DCFM-B2 and DCFM-B5 improve FPS from 44.1 and 13.9 to 67.6 and 25.4 (speed up 1.5 times and 1.8 times), respectively, while maintaining approximately 1% higher mIoU with the keyframe interval $K$ = 2. In comparison to MRCFA-B5, the current state-of-the-art method on VSPW in terms of accuracy, DCFM-B5 ($K$ = 2) exhibits a slight decrease of -0.6% in mIoU, but it achieves around fivefold increase in FPS (from 5.3 to 25.4). By adjusting $K$ to 6 during inference, our DCFM-B5 model achieves a significantly faster speed. Despite the slight decrease in mIoU, it still outperforms MRCFA-B2 and CFFM-B2 in terms of both accuracy and inference speed, despite having a larger parameter count. The results visualized in Figure 6a further demonstrate superior temporal consistency (mVC) and a more favorable trade-off between accuracy and efficiency achieved by the proposed approach with fewer parameters. Furthermore, our model tends to be easier to train due to its simple architecture without feature recalibration module as well as the proposed symmetric training scheme.
TABLE III: Performance comparison in terms of accuracy (mIoU) and inference speed (FPS) on the Cityscapes [26] validation dataset. Methods, excluding ours, are arranged based on mIoU. DCFM* indicates a variant of our model with a different input resolution ($640\times 1280$). Results are visualized in Figure 6b.
Methods Backbone mIoU(%) $\uparrow$ FPS $\uparrow$
@@ IV-C Comparison with previous approaches
TABLE IV: Performance comparison in terms of accuracy (mIoU) and inference speed (FPS) on the CamVid [69] test dataset. Methods, excluding ours, are arranged based on mIoU.
Methods Backbone mIoU(%) $\uparrow$ FPS $\uparrow$
DFF [63] ResNet-101 66.0 16.1
GRFP [58] ResNet-101 66.1 4.3
Accel [56] ResNet-18 66.7 7.6
TDNet [59] ResNet-18 72.6 25.0
DCNet [10] ResNet-50 73.0 20.0
DCFM (Ours) ResNet-50 73.1 23.7
(a) VSPW
(b) Cityscapes
Figure 6: Accuracy (mIoU) and inference speed (FPS) comparison on VSPW [14] and Cityscapes [26]. In the figure, the yellow square, green triangle, and blue circle represent methods using ResNet [1], Swin [71], and MiT [2] as the backbone network, respectively. Our method is indicated by a red border.
@@
We extend our model comparison to the Cityscapes [26], detailed in Table III and Figure 6b. DCFM-50 outperforms contemporary feature aggregation methods like TDNet, which prioritizes high accuracy while utilizing multiple distributed networks for enhanced inference speed simultaneously. Notably, our method achieves a superior mIoU while maintaining a higher inference speed. In comparison to the method with the highest mIoU score, LMA [17], our DCFM-101 exhibits a 0.5% lower mIoU but is 4.3 times faster. When compared with other keyframe-based techniques aimed at accelerating segmentation models, we surpass LVS [9] and DCNet [10] in both accuracy (77.2% mIoU) and speed (13.2 FPS) with a smaller backbone (ResNet-50 [1]). These two methods also leverage different layers of the backbone network for feature reuse like ours, while our approach stands out by decoupling keyframe information into the common and independent parts. Among the methods using MiT [2] as the backbone, DCFM is both faster and more accurate than the SegFormer [2] baseline. Compared with CFFM [15] and MRCFA [15], the most accurate methods on VSS, DCFM achieves 0.7% less mIoU but about twice as fast.
Additionally, we validate the effectiveness of DCFM on the CamVid [69] dataset. As shown in the Table IV, DCFM achieves the highest accuracy while its FPS is only 1.3 less than the fastest method, TDNet [59]. Additionally, DCFM outperforms the latest keyframe-based method, DCNet [10], in both mIoU and FPS.
The results on these three datasets confirm the effectiveness of deep common feature mining, a methodology fundamentally distinct from feature aggregation over multiple frames (as shown in Figure 2b) and keyframe-based feature propagation (as shown in Figure 2c). In contrast to the former, mining common features across neighboring frames significantly reduces redundant computation. When compared to the latter, the decomposition operation in our model eliminates the need for unreliable feature re-calibration. Further insights into the advantage of feature decoupling are presented in the ablation experiments detailed in Section IV-D.
TABLE V: Quantitative ablation study on the decoupled feature representations, achieved on the VSPW [14] validation dataset. mIoUk and mIoUn indicate the mIoU metric calculated on keyframes and non-key frames, respectively.
@@
44.2 43.7 44.0
3.9 42.7 24.3
44.2 43.4 43.8
43.8 3.59 24.7
45.0 44.6 44.8
TABLE VI: Quantitative ablation study on the loss functions, achieved by DCFM-B2 on the VSPW [28] validation dataset.
Loss function mIoU (%) $\text{mVC}_{8}$(%) $\text{mVC}_{16}$(%)
$L_{I}$ $L_{B}$ $L_{C}$
43.5 87.4 82.5
43.0 87.8 83.3
43.8 87.7 83.0
45.0 88.1 83.5
TABLE VII: Quantitative ablation study using different functions as the loss function for DCFM, achieved by DCFM-B2 on the VSPW [14] validation dataset.
Loss function mIoU (%) $\text{mVC}_{8}$(%) $\text{mVC}_{16}$(%)
$L_{I}$&$L_{B}$ $L_{C}$
CE+Dice [73] L1 42.9 88.1 83.4
CE+Dice [73] MSE 44.2 88.6 84.1
OhemCE [74] L1 44.5 88.0 83.4
OhemCE [74] MSE 43.7 88.2 83.7
CE L1 43.5 87.4 82.6
CE MSE 45.0 88.1 83.5
IV-D Ablation study
In this section, we experimentally verify the efficacy of several designs incorporated into our model.
Effectiveness of feature decomposing. The core idea of our method is to mine the common representation between frames and reuse it during inference directly. By decomposing features into common and independent representations, our approach distinguishes itself from other keyframe-based methods. In Table V, we perform ablation study by omitting one of the feature components in keyframes or non-key frames during both training and testing, where MiT-B2 is exploited as the backbone. As shown in the first row, excluding the shallow features $F^{key}_{in}$ of the keyframe causes our framework to revert to the structure of previous feature propagation-based methods, as depicted in Figure 2b. The observed significant decrease in accuracy suggests a misalignment between the keyframes and the corresponding non-keyframes. However, instead of applying a re-calibration operation on the utilized feature to address this issue, our framework mitigates misalignments by mining common information across frames. In the experiment of the second row, the model structure on non-key frames remains unchanged, but the accuracy mIoUn still decreases significantly. This suggests that the mining of common features and the proposed symmetric training strategy are essential. Comparing the results of the ablation model in the third row with those obtained by the full model, it is evident that the learned common feature is highly informative, containing most of the necessary information for the non-key frames. However, utilizing the common feature $F^{key}_{co}$ provided by the keyframe alone on non-key frames is suboptimal. This is primarily because it neglects the quickly changing frame-specific dynamics, and such complementary information is well captured by the independent feature in our full model. The third and fourth lines in the table show that using one of these two parts alone on non-key frames is not feasible. This suggests that the exceptional performance of our method on non-key frames is not solely due to the learnable network parameters but also relates to feature decomposition.
(a) w/o $L_{C}$
(b) w/ $L_{C}$
Figure 8: t-SNE [72] visualization of the features extracted by the models trained without (a) or with (b) $L_{C}$. Points in different colors represent feature vectors from distinct semantic categories.
Figure 9: Results of adaptive keyframe scheduling on VSPW [14].
TABLE VIII: Comparison of the inference speed on VSPW [14].
@@ IV-D Ablation study
Effectiveness of loss functions. Ablation studies on all training losses, as shown in Table VI, validate their effectiveness. Using only $L_{I}$ results in the DCFM model behaving like a standard image semantic segmentation model, with unsatisfactory performance due to inadequate consideration of temporal correlations, as shown in the first row. Solely employing $L_{B}$ ensures consistent utilization of deep features from adjacent frames for segmentation during training. Despite a slight decrease in accuracy during testing, there is an improvement in temporal consistency, with the $\text{mVC}_{16}$ value rising from 82.5% to 83.3%. Joint utilization of $L_{I}$ and $L_{B}$ improves mIoU accuracy from 43.5% to 43.8%. Results in the fourth row illustrate that including $L_{C}$ significantly enhances segmentation accuracy and temporal consistency, particularly in terms of consistency over long periods of time ($\text{mVC}_{16}$). Figure 7 visually evaluates the impact by calculating the cosine similarity between each pixel and its surrounding eight pixels in the feature space (calculated on the output of FFM), indicating that the learned features are less susceptible to rapidly changing details in areas with the same semantic category when $L_{C}$ is applied. Furthermore, we randomly select 300 feature vectors for each category when testing DCFM on Cityscapes [26] and visualize them using t-SNE [72]. As depicted in Figure 8, the consistency loss encourages the model to learn a more robust feature representation with lower intra-class dispersion and reduced errors.
we have also conducted ablation experiments to explore the impact of alternative loss functions such as Dice loss [73], OhemCE [74], and L1 loss. These experiments, detailed in Table VII , reveal that while Dice loss and OhemCE lead to a slight reduction (approximately 1%) in segmentation accuracy (mIoU), they provide minor improvements for $\text{mVC}_{8}$ by 0.5% and 0.1%. Conversely, L1 loss consistently performs worse than MSE in our context. These results confirm that while our framework is robust to different loss functions, MSE and cross-entropy remain effective choices for our specific task, contributing to the overall performance of the proposed method.
@@
Analysis on inference speed. We experimentally analyze the inference time of the modules in DCFM-B5 on the VSPW [14] dataset, as shown in Table VIII. The average inference time of DCFM is $\frac{(K-1)T_{n}+T_{k}}{k}$, where $T_{k}$ and $T_{n}$ respectively indicate the inference time on keyframes and non-key frames. For the keyframe, as DCFM only uses an additional lightweight FFM module to integrate features before the segmentation head, the inference time increases by only 1.1ms compared with the SegFormer-B5 [2] baseline, accounting for 1.5% of the total time spent on a keyframe. Regarding the non-keyframes, the computing capacity of the adopted shallow network has only 1.48 GFLOPS, thus DCFM achieves a very fast speed of 5.8ms (172 FPS) by reusing the keyframe common features, constituting only 8% of the keyframe inference time. As keyframe-based semantic segmentation methods, DCNet [10] and LVS [9] respectively occupy 29% and 33% of the inference time on non-keyframes compared to keyframes. Therefore, DCFM exhibits a considerable advantage in inference speed on non-keyframes, implying that in certain fixed scenes or high frame-rate videos, DCFM can significantly reduce the required inference time by increasing the interval between keyframes while maintaining segmentation accuracy.
Additionally, as demonstrated in Table IX, we have evaluated the performance of the proposed method when re-utilizing only previous keyframes (“P mode”) in comparison to employing both previous and future keyframes (“B mode”) for segmentation. The results indicate that bidirectional dependency or reliance solely on preceding keyframes does not significantly impact the overall running speed of DCFM. This efficiency is primarily attributed to the effective reuse of deep features, which incurs minimal additional computational overhead. Moreover, the flexibility to select the testing mode according to specific application scenarios further enhances the adaptability of the method.
(a) VSPW
(b) Cityscapes
Figure 10: Result visualization of VSS methods on VSPW [14] and Cityscapes [26]. (a) Comparison of segmentation results on the VSPW [14] dataset with SegFormer [2] and CFFM [15]. The examples in the figure include three frames from each of two videos, demonstrating DCFM’s advantages in background consistency and high-speed moving object segmentation. For DCFM, the first and third columns are segmentation results when the images are treated as non-key frames. (b) Comparison of segmentation results on the Cityscapes dataset with DCNet. The examples in the figure include three annotated frames from three different videos. The columns labeled “Key” show the segmentation results of the two models when the frames are used as keyframes, while the columns labeled “Kon-key” show the segmentation results when the frames are used as non-key frames, with a same distance of 1 from the nearest keyframe.
Analysis on adaptive keyframe scheduling. As depicted in Figure 9, the accuracy and speed of DCFM-B5 exhibit consistent changes with the increase of the keyframe interval $K$. This figure also presents the results of DCFM-B5 with adaptive keyframe scheduling (AKS). For comparison, we set the maximum value of fixed interval value $K$ to be 6. For AKS, the minimum value MinK is set to 1 or 2. By adjusting the value of the threshold $S$, AKS significantly achieves a higher accuracy under a same average keyframe interval on DCFM-B5, indicating that our model performance can be further improved without re-training. The results also suggest that our method and the adaptive keyframe scheduling are orthogonal, allowing for the selection of specific algorithms in practical applications. In comparison with other VSS methods, AKS is excluded from the proposed DCFM. For methods such as DCNet [10], LVS [9], and DFF [63], we consider their best results with AKS for comparison, thereby demonstrating the effectiveness and efficiency of our approach.
V Result visualization
We have performed a visual comparison of the video segmentation results on the frame-by-frame annotated VSPW [14] dataset with the image segmentation baseline SegFormer [2] and CFFM [15], as shown in Figure 10a. The background in the first video clearly demonstrates that DCFM outperforms SegFormer and CFFM in both spatial and temporal consistency. Observing the helicopter rotor in the second video, DCFM demonstrates superior performance in handling high-speed moving objects compared to the other methods shown. Additionally, it exhibits a significant advantage in temporal consistency for segmenting distant buildings.
Additionally, we perform a comparative evaluation between the keyframe-based representative method DCNet [10] and the proposed DCFM using the same experimental setup on the Cityscapes dataset [26]. The segmentation results are visually illustrated in Figure 10b. Notably, our method demonstrates significantly improved consistency with the input frames compared to DCNet. This observation suggests that the propagation module in DCNet tends to carry unnecessary information over to non-key frames, leading to an excessive similarity between segmentation results of key and non-key frames. Specifically, the segmentation errors on non-key frames are evident, as exemplified by buses and vegetation within the white box in the figure.
VI Conclusion
TABLE IX: Quantitative ablation study on the inference mode of non-key frames, achieved by DCFM-B2 and DCFM-b5 on the VSPW [28] validation dataset. $P$ represents a non-key frame that only relies on the nearest previous keyframe for segmentation, while $B$ represents a non-key frame that relies on both the nearest previous and future key frames for segmentation.
Methods Mode mIoU(%) Speed (ms/f) Latency (ms)
DCFM-B2 $B$ 44.8 14.8 90.5
$P$ 44.5 13.8 23.8
DCFM-B5 $B$ 49.3 39.4 136.7
$P$ 48.9 38.4 70.0
@@ VI Conclusion
This study addresses challenges in efficient video semantic segmentation. The proposed DCFM method aims to maintain high accuracy and temporal consistency while overcoming computational bottlenecks and feature propagation issues in existing approaches. It decomposes intricate features into two components, i.e., common and independent representations, where the former can be directly re-utilized by neighboring frames without additional re-calibration, contributing to an optimized trade-off between accuracy and speed. A symmetric training strategy and a self-supervised consistency loss are introduced, enhancing the ability to handle sparsely annotated data and offering potential applications in other video analysis tasks.
In contrast to existing keyframe-based VSS methods, DCFM leverages common feature mining instead of propagation modules such as optical flow, thereby enhancing computational efficiency. This efficiency is attained through the effective reuse of deep features, resulting in minimal additional computational overhead. Notably, the inference speed for non-keyframes in our method is only 8% of that for keyframes. This design renders DCFM particularly well-suited for applications in scenarios such as security surveillance and high-resolution, high-frame-rate contexts like live video streaming, where the benefits of keyframe feature reuse can be maximized. However, similar to many keyframe-based VSS methods, the primary latency bottleneck is associated with the segmentation of keyframes, resulting in higher overall latency compared to the average segmentation time per frame. While methods like LVS [9] reduce latency by offloading the high-precision segmentation of keyframes to the background, real-time image semantic segmentation methods (e.g., STDC [55], BiSeNet [50]) are significantly more suitable for scenarios requiring low latency or low computational power [75, 76], such as autonomous driving or robotics, due to their advantages in latency and hardware optimization over VSS methods. Future work will focus on further optimizing DCFM for real-time applications and exploring additional strategies to reduce latency.
Acknowledgments
This work is partly supported by the National Key R&D Program of China (2022ZD0161902), the Research Program of State Key Laboratory of Software Development Environment, and the Fundamental Research Funds for the Central Universities.